# Torsion and Chern-Simons gravity in 4D space-times from a Geometrodynamical four-form

### v2

@@ Abstract
The above equation can be solved exactly and, out of the four roots, only one is real and positive. However, given the smallness of the quantity $\lambda_{M}\equiv\Lambda c^{2}/6GM\cong 10^{-68}\ cm^{-3}$ in our case, even up to the cubic expansion of this root, $R_{bec}$, in terms of $\lambda_{M}$, the solution to eq.(51) corresponding to the minimum energy configuration of the BEC is (upto $\lambda^{3}_{M}$ orders),
$R_{bec}\cong\frac{\hbar^{2}}{4GMm^{2}}\cong 22\bigg{(}\frac{10^{9}M_{\odot}}{M}\bigg{)}\bigg{(}\frac{10^{-22}\ \mbox{eV}}{m}\bigg{)}^{2}\ \mbox{pc}$ (52),52( 52 ) ,
corresponding to a single boson energy,
$E_{min}=E_{b}(R_{bec})\cong-\frac{GMm}{2R_{bec}}=-2\ mc^{2}\bigg{(}\frac{m^{2}_{Pl}}{m\ M}\bigg{)}^{-2}$ (53).53( 53 ) .
@@
Eqs.(52) and (53) are exact when $\Lambda=0$, ensuing trivially from eq.(51). These equations entail that for larger BEC mass $M$ not only its size is smaller, it is also more tightly bound since its energy is more negative. The physical repercussion of these features is that for a sufficiently large mass $M$, $R_{bec}$ can be very close to the corresponding Schwarzschild radius $R_{s}\equiv 2GM/c^{2}$ so that even a slight perturbation would cause an irreversible gravitational collapse leading to the formation of a BH.
@@
The BEC would form a BH if its size $R_{bec}$ goes below the Schwarzschild radius $R_{S}$,
$R_{S}=\frac{2GM}{c^{2}}=\frac{2M}{m^{2}_{Pl}}$ (54),54( 54 ) ,
where $m_{Pl}$ is the Planck mass.
@@
Using the criteria $R_{bec}\lesssim R_{S}$ along with eqs.(52) to (54), the condition for the BEC to implode into a BH is given by the inequality,
$\bigg{(}\frac{m^{2}_{Pl}}{4M\ m}\bigg{)}^{2}\lesssim 1$ (55),55( 55 ) ,
$\Rightarrow\ m\ M\gtrsim\ 0.25\ m^{2}_{Pl}=3\times 10^{9}\ \bigg{(}\frac{m}{10^{-20}\ \mbox{eV}}\bigg{)}^{-1}M_{\odot}$ (56).56( 56 ) .
@@
A more rigorous analysis carried out earlier by employing a Gross-Pitaevskii equation framework had shown that the dynamical evolution of ultra-light dark bosons in the BEC phase leads to the formation of a BH on time scales of $10^{8}$ years with a similar constraint given by eq.(56) PDGEklavya ; PDGFazlu .
The Hawking temperature $T_{BH}$ of a BH of mass $M$ is given by Hawking ,
$k_{B}T_{BH}=\frac{m^{2}_{Pl}}{8\pi M}$ (57).57( 57 ) .
It is interesting to note that when eq.(56) is substituted in eq.(57), one gets,
$k_{B}T_{BH}\lesssim\frac{m}{2\pi}$ (59),59( 59 ) ,
@@
implying that a SMBH created from the collapse of a BEC of dark bosons would be predominantly radiating massless particles as well as these dark bosons through Hawking evaporation!
V V. Propagation of Gravitational Waves in the presence of the geometrodynamical four-form
@@
We consider in this section the effect of $\tilde{w}$ on the propagation of gravitational waves (GWs) by assuming that, except for a weak GW and the geometrodynamical four form, there is no other matter present. Considering a traceless-transverse (TT)-gauge for a plane GW travelling along the x-axis,
$h_{\oplus}\equiv h_{22}(t,x)=-h_{33}(t,x)\ \ \ \ \ h_{\otimes}\equiv h_{23}(t,x)=h_{32}(t,x)$
@@
we express eqs.(19) and (28), by making the weak field approximation $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ and retaining only upto quadratic terms in $h_{\oplus}$ and $h_{\otimes}$ for the gravitational part of the action Alex ,
$\mathcal{A}\approx-{{m^{2}_{Pl}}\over{8\pi}}\int{[h_{\oplus}(\ddot{h}_{\oplus}-h^{\prime\prime}_{\oplus})+h_{\otimes}(\ddot{h}_{\otimes}-h^{\prime\prime}_{\otimes})]d^{4}x}+\mathcal{A}_{GCS}+\mathcal{A}_{\phi}$ (60)60( 60 )
@@
$\ \ \ \ \ \ \ \ \ +H\int{\frac{\partial\ln\chi}{\partial t}\bigg{(}h_{\oplus}(\ddot{h}_{\otimes,1}-h^{\prime\prime}_{\otimes,1})-h_{\otimes}(\ddot{h}_{\oplus,1}-h^{\prime\prime}_{\oplus,1})\bigg{)}d^{4}x}$
@@
$\mathcal{A}_{\phi}=\int{\bigg{[}\frac{a_{1}}{\chi}\eta^{\mu\nu}\chi_{,\mu}\chi_{,\nu}+a_{2}\chi-\frac{a_{1}}{\chi}\bigg{(}h_{\oplus}(\chi_{,2}\chi_{,2}-\chi_{,3}\chi_{,3})+2h_{\otimes}\chi_{,2}\chi_{,3}\bigg{)}\bigg{]}d^{4}x}$
After defining $\psi(x^{\mu})\equiv\ln{(\chi(x^{\mu}))}$, the dynamical equations of motion that ensue from variations of $\mathcal{A}$ (eq.(60)) with respect to $h_{\oplus}$, $h_{\otimes}$ and $\psi$ are given by,
@@
$\ \ \ \ \ \ \ \ \ +a_{1}\exp{(\psi)}\bigg{\{}\bigg{(}\frac{\partial\psi}{\partial y}\bigg{)}^{2}-\bigg{(}\frac{\partial\psi}{\partial z}\bigg{)}^{2}\bigg{\}}\bigg{]}$ (61)61( 61 )
@@
$\ \ \ \ \ \ \ \ \ \ +2a_{1}\exp{(\psi)}\frac{\partial\psi}{\partial y}\frac{\partial\psi}{\partial z}\bigg{]}$ (62)62( 62 )
@@
$\partial^{\mu}\partial_{\mu}\psi+\frac{1}{2}\eta^{\mu\nu}\frac{\partial\psi}{\partial x^{\mu}}\frac{\partial\psi}{\partial x^{\nu}}-\frac{a_{2}}{4a_{1}}=h_{\oplus}\bigg{[}\frac{\partial^{2}\psi}{\partial y^{2}}-\frac{\partial^{2}\psi}{\partial z^{2}}+\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial y}\bigg{)}^{2}-\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial z}\bigg{)}^{2}\bigg{]}+2h_{\otimes}\bigg{[}\frac{\partial^{2}\psi}{\partial y\partial z}+\frac{1}{2}\frac{\partial\psi}{\partial y}\frac{\partial\psi}{\partial z}\bigg{]}\ \ .$ (63)63( 63 )
@@
A simplification occurs if we assume $\psi$ to depend only on the x-coordinate and time, like the GW amplitude itself. With $\psi=\psi(t,x)$, the last terms in eqs.(61) and (62) drop out. Introducing the complex circularly polarized GW amplitudes [21],
$h_{R}=\frac{1}{\sqrt{2}}(h_{\oplus}+ih_{\otimes})\ \ \ \ \ \ h_{L}=\frac{1}{\sqrt{2}}(h_{\oplus}-ih_{\otimes})$
one finds from eqs.(61) and (62) that,
$\ddot{h}_{R}-h^{\prime\prime}_{R}=-\frac{8\pi Hi}{m^{2}_{Pl}}\bigg{[}\dot{\psi}\frac{\partial}{\partial x}(\ddot{h}_{R}-h^{\prime\prime}_{R})-\psi^{\prime}\frac{\partial}{\partial t}(\ddot{h}_{R}-h^{\prime\prime}_{R})+\dot{h}^{\prime}_{R}(\ddot{\psi}+\psi^{\prime\prime})-\dot{\psi}^{\prime}(\ddot{h}_{R}+h^{\prime\prime}_{R})\bigg{]}$ (64)64( 64 )
$\ddot{h}_{L}-h^{\prime\prime}_{L}=\frac{8\pi Hi}{m^{2}_{Pl}}\bigg{[}\dot{\psi}\frac{\partial}{\partial x}(\ddot{h}_{L}-h^{\prime\prime}_{L})-\psi^{\prime}\frac{\partial}{\partial t}(\ddot{h}_{L}-h^{\prime\prime}_{L})+\dot{h}^{\prime}_{L}(\ddot{\psi}+\psi^{\prime\prime})-\dot{\psi}^{\prime}(\ddot{h}_{L}+h^{\prime\prime}_{L})\bigg{]}$ (65)65( 65 )
Since $\psi=\psi(t,x)$, the equation of motion for the four-form given by eq.(63) simplifies to,
@@
where $\mu\equiv\sqrt{-\frac{a_{2}}{4a_{1}}}$ acts as the mass of the four-form field. Seeking exact solutions, involving monochromatic and circularly polarized GWs, we substitute,
@@
in eq.(64), leading to a relation,
$\omega^{2}-k^{2}=-\beta[k(\omega^{2}-k^{2})\dot{\psi}+\omega(\omega^{2}-k^{2})\psi^{\prime}-i\omega k(\ddot{\psi}+\psi^{\prime\prime})-i(\omega^{2}+k^{2})\dot{\psi}^{\prime}]$ (67)67( 67 )
@@
If $w=\pm k$, eq.(67) implies,
$2\dot{\psi}^{\prime}\pm(\ddot{\psi}+\psi^{\prime\prime})=0\ \ .$ (68)68( 68 )
The only self-consistent solution of eqs.(66) and (68) is,
$\psi(t,x)=(b_{0}-\frac{\mu^{2}}{b_{0}})t+(b_{0}+\frac{\mu^{2}}{b_{0}})x$ (69)69( 69 )
where $b_{0}$ is an integration constant. But the above solution is unphysical, as it implies an exponentially growing $\psi$. On the other hand, when $w^{2}>k^{2}$,
$k\dot{\psi}+\omega\psi^{\prime}+\frac{1}{\beta}=0\Rightarrow\psi^{\prime\prime}=\frac{k^{2}}{w^{2}}\ \ddot{\psi}$ (70)70( 70 )
@@
$\psi=\ln{\bigg{(}\sqrt{\frac{b_{1}}{\lambda}}\cos^{2}\bigg{(}\pm\sqrt{\frac{\lambda}{2}}(t-\frac{k}{\omega}x)+b_{2}\bigg{)}\bigg{)}}$ (71)71( 71 )
@@
Hence, $\chi(t-\frac{k}{\omega}x)=\exp(\psi)\propto\cos^{2}\bigg{(}\pm\sqrt{\frac{\lambda}{2}}(t-\frac{k}{\omega}x)+b_{2}\bigg{)}$, with $|\omega|>|k|$, is physically meaningful and is an acceptable solution. Indeed, we find that it is possible to have circularly polarized, monochromatic GWs with phase velocity exceeding the speed of light. The caveat, however, is that eq.(66) being a nonlinear differential equation, deriving exact solutions corresponding to GWs with superposed wave modes is non-trivial.
VI VI. Conclusions
Based on the invariances of the Minkowski tensor and the flat space-time Levi-Civita tensor under proper Lorentz transformations, the present study explores the possibility of extending Einstein’s geometrical theory of gravitation by including another geometrical degree of freedom $\tilde{w}$, which is simply a generalisation of the Levi-Civita symbol, in the theory. This geometrodynamical four-form $\tilde{w}$ leads not only to a dynamical torsion it also generates Chern-Simon (CS) extensions in the 3+1-dimensional space-times. Torsion, among several other interesting implications, is also important in teleparallel gravity TelepGravity .
Pursuing closely the seminal work of Jackiw and Pi Jackiw but avoiding the use of an unphysical Lorentz vector, CS coupling between electromagnetic fields and $\tilde{w}$ comes about naturally, bringing about a modification of Einstein-Maxwell equations. Adopting once again the formulation of Jackiw , a gravitational CS term is constructed that leads to a modified Cotton tensor. However, when the dynamical equation of the four-form is used in the ensuing Einstein equation, it is shown that it is the standard Cotton tensor that affects the space-time dynamics.

### v3

@@ 4 Dark bosons, Bose-Einstein Condensates and Formation of Supermassive Black Holes
The above equation can be solved exactly and, out of the four roots, only one happens to be real and positive. However, given the smallness of the quantity $\lambda_{M}$, this real and positive solution to eq.(4.11), even up to $\lambda^{3}_{M}$ orders, is of the form,
$R_{bec}\cong\frac{\hbar^{2}}{4GMm^{2}}\cong 22\bigg{(}\frac{10^{9}M_{\odot}}{M}\bigg{)}\bigg{(}\frac{10^{-22}\ \mbox{eV}}{m}\bigg{)}^{2}\ \mbox{pc}\ \ ,$ (4.12)
that corresponds to a single boson energy,
$E_{min}=E_{b}(R_{bec})\cong-\frac{GMm}{2R_{bec}}=-2\ mc^{2}\bigg{(}\frac{m^{2}_{Pl}}{m\ M}\bigg{)}^{-2}\ .$ (4.13)
Eqs.(4.12) and (4.13) are of course exact solutions for the $\Lambda=0$ case, ensuing trivially from eq.(4.11).
@@
The preceding equations entail that for larger BEC mass $M$ or larger rest mass $m$ of the pseudo-scalar not only the BEC size is smaller, the condensate is also more tightly bound since its energy is more negative. The physical repercussion of these features is that for a sufficiently large mass $M$ or $m$, $R_{bec}$ can be very close to the corresponding Schwarzschild radius $R_{s}\equiv 2GM/c^{2}$ so that even a slight perturbation would cause an irreversible gravitational collapse leading to the formation of a BH.
@@
The BEC would form a BH if its size $R_{bec}$ shrinks below the Schwarzschild radius $R_{S}$ limit,
$R_{S}=\frac{2GM}{c^{2}}=\frac{2M}{m^{2}_{Pl}}\ \ ,$ (4.14)
where $m_{Pl}$ represents the Planck mass.
@@
Using the criteria $R_{bec}\lesssim R_{S}$ along with eqs.(4.12) and (4.14), the condition for the BEC to implode into a BH is given by the inequality,
$\bigg{(}\frac{m^{2}_{Pl}}{4M\ m}\bigg{)}^{2}\lesssim 1\ ,$ (4.15)
$\Rightarrow\ m\ M\gtrsim\ 0.25\ m^{2}_{Pl}=3\times 10^{9}\ \bigg{(}\frac{m}{10^{-20}\ \mbox{eV}}\bigg{)}^{-1}M_{\odot}\ .$ (4.16)
@@
A more rigorous analysis carried out earlier, for the $\Lambda=0$ case, by employing a Gross-Pitaevskii equation framework had shown that the dynamical evolution of ultra-light dark bosons in the BEC phase leads to the formation of a BH on time scales $\sim 10^{8}$ years with a similar constraint given by eq.(4.16) [17, 18].
Now, the Hawking temperature $T_{BH}$ of a BH of mass $M$ is given by [20],
$k_{B}T_{BH}=\frac{m^{2}_{Pl}}{8\pi M}\ \ .$ (4.17)
So, it is interesting to note that when eq.(4.16) is substituted in eq.(4.17), one gets,
$k_{B}T_{BH}\lesssim\frac{m}{2\pi}\ \ .$ (4.18)
@@
In other words, a SMBH created from the collapse of a BEC of ultra-light dark bosons would be predominantly radiating, via Hawking evaporation, massless particles as well as these pseudo-scalars associated with $\tilde{w}$.
5 Propagation of Gravitational Waves in the presence of the geometrodynamical four-form
@@
We take up in this section the effect of $\tilde{w}$ on the propagation of gravitational waves (GWs) assuming that, except for a weak GW and the geometrodynamical four form, there is no other matter present.
By making the weak field approximation $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ (with $|h_{\mu\nu}|\ll 1)$) and considering a traceless-transverse (TT)-gauge for a plane GW travelling along the x-axis, we express the GW amplitudes,
$h_{\oplus}\equiv h_{22}(t,x)=-h_{33}(t,x)\ \ \ \mbox{and}\ \ \ h_{\otimes}\equiv h_{23}(t,x)=h_{32}(t,x)\ .$ (5.1)
@@
Retaining only upto quadratic terms in $h_{\oplus}$ and $h_{\otimes}$ for the gravitational part, the action given by eqs.(3.5) and (3.14) reduce to (also, see [21]),
$\mathcal{A}\approx-{{m^{2}_{Pl}}\over{8\pi}}\int{[h_{\oplus}(\ddot{h}_{\oplus}-h^{\prime\prime}_{\oplus})+h_{\otimes}(\ddot{h}_{\otimes}-h^{\prime\prime}_{\otimes})]d^{4}x}+\mathcal{A}_{GCS}+\mathcal{A}_{\phi}$ (5.2)
@@
$\ \ \ \ \ \ \ \ \ +H\int{\frac{\partial\ln\chi}{\partial t}\bigg{(}h_{\oplus}(\ddot{h}_{\otimes,1}-h^{\prime\prime}_{\otimes,1})-h_{\otimes}(\ddot{h}_{\oplus,1}-h^{\prime\prime}_{\oplus,1})\bigg{)}d^{4}x}$ (5.3)
@@
$\mathcal{A}_{\phi}=\int{\bigg{[}\frac{a_{1}}{\chi}\eta^{\mu\nu}\chi_{,\mu}\chi_{,\nu}+a_{2}\chi-\frac{a_{1}}{\chi}\bigg{(}h_{\oplus}(\chi_{,2}\chi_{,2}-\chi_{,3}\chi_{,3})+2h_{\otimes}\chi_{,2}\chi_{,3}\bigg{)}\bigg{]}d^{4}x}\ \ .$ (5.4)
After defining $\psi(x^{\mu})\equiv\ln{(\chi(x^{\mu}))}$, the dynamical equations of motion that ensue from variations of $\mathcal{A}$ (eq.(5.2)) with respect to $h_{\oplus}$, $h_{\otimes}$ and $\psi$ are given by,
@@
$\ \ \ \ \ \ \ \ \ +a_{1}\exp{(\psi)}\bigg{\{}\bigg{(}\frac{\partial\psi}{\partial y}\bigg{)}^{2}-\bigg{(}\frac{\partial\psi}{\partial z}\bigg{)}^{2}\bigg{\}}\bigg{]}\ ,$ (5.5)
@@
$\ \ \ \ \ \ \ \ \ \ +2a_{1}\exp{(\psi)}\frac{\partial\psi}{\partial y}\frac{\partial\psi}{\partial z}\bigg{]}$ (5.6)
@@
$\partial^{\mu}\partial_{\mu}\psi+\frac{1}{2}\eta^{\mu\nu}\frac{\partial\psi}{\partial x^{\mu}}\frac{\partial\psi}{\partial x^{\nu}}-\frac{a_{2}}{4a_{1}}=h_{\oplus}\bigg{[}\frac{\partial^{2}\psi}{\partial y^{2}}-\frac{\partial^{2}\psi}{\partial z^{2}}+\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial y}\bigg{)}^{2}-\frac{1}{2}\bigg{(}\frac{\partial\psi}{\partial z}\bigg{)}^{2}\bigg{]}+2h_{\otimes}\bigg{[}\frac{\partial^{2}\psi}{\partial y\partial z}+\frac{1}{2}\frac{\partial\psi}{\partial y}\frac{\partial\psi}{\partial z}\bigg{]}\ \ .$ (5.7)
@@
A simplification occurs if we assume $\psi$ to depend only on the x-coordinate and time, like the GW amplitude itself. With $\psi=\psi(t,x)$, the last terms in eqs.(5.5) and (5.6) drop out. Introducing the circularly polarized GW amplitudes, as demonstrated in the reference [21],
$h_{R}=\frac{1}{\sqrt{2}}(h_{\oplus}+ih_{\otimes})\ \ \ \mbox{and}\ \ \ h_{L}=\frac{1}{\sqrt{2}}(h_{\oplus}-ih_{\otimes})\ ,$
one finds that the coupled differential equations given by eqs.(5.5) and (5.6), involving $h_{\oplus}$ and $h_{\otimes}$, get mutually separated,
$\ddot{h}_{R}-h^{\prime\prime}_{R}=-\frac{8\pi Hi}{m^{2}_{Pl}}\bigg{[}\dot{\psi}\frac{\partial}{\partial x}(\ddot{h}_{R}-h^{\prime\prime}_{R})-\psi^{\prime}\frac{\partial}{\partial t}(\ddot{h}_{R}-h^{\prime\prime}_{R})+\dot{h}^{\prime}_{R}(\ddot{\psi}+\psi^{\prime\prime})-\dot{\psi}^{\prime}(\ddot{h}_{R}+h^{\prime\prime}_{R})\bigg{]}$ (5.8)
and
$\ddot{h}_{L}-h^{\prime\prime}_{L}=\frac{8\pi Hi}{m^{2}_{Pl}}\bigg{[}\dot{\psi}\frac{\partial}{\partial x}(\ddot{h}_{L}-h^{\prime\prime}_{L})-\psi^{\prime}\frac{\partial}{\partial t}(\ddot{h}_{L}-h^{\prime\prime}_{L})+\dot{h}^{\prime}_{L}(\ddot{\psi}+\psi^{\prime\prime})-\dot{\psi}^{\prime}(\ddot{h}_{L}+h^{\prime\prime}_{L})\bigg{]}\ ,$ (5.9)
displaying a symmetry under $R\longleftrightarrow L$ in the above equations modulo a negative sign, corresponding to the two independent circular polarizations.
Since $\psi=\psi(t,x)$, the equation of motion for the four-form given by eq.(5.7) simplifies to,
@@
where $\mu\equiv\sqrt{-\frac{a_{2}}{4a_{1}}}$ is akin to the rest mass of the four-form field.
Seeking exact solutions of eqs.(5.8) and (5.10), involving monochromatic and circularly polarized GWs as well as $\psi$, we substitute,
@@
in eq.(5.8), leading to a relation,
$\omega^{2}-k^{2}=-\beta[k(\omega^{2}-k^{2})\dot{\psi}+\omega(\omega^{2}-k^{2})\psi^{\prime}-i\omega k(\ddot{\psi}+\psi^{\prime\prime})-i(\omega^{2}+k^{2})\dot{\psi}^{\prime}]$ (5.11)
@@
If $w=\pm k$, eq.(5.11) implies,
$2\dot{\psi}^{\prime}\pm(\ddot{\psi}+\psi^{\prime\prime})=0\ \ .$ (5.12)
The only self-consistent solution of eqs.(5.10) and (5.12) is,
$\psi(t,x)=(b_{0}-\frac{\mu^{2}}{b_{0}})t+(b_{0}+\frac{\mu^{2}}{b_{0}})x$ (5.13)
where $b_{0}$ is an integration constant. But the above solution is unphysical, as it implies an exponentially growing $\chi(t,x)$. On the other hand, when $w^{2}>k^{2}$,
$k\dot{\psi}+\omega\psi^{\prime}+\frac{1}{\beta}=0\Rightarrow\psi^{\prime\prime}=\frac{k^{2}}{w^{2}}\ \ddot{\psi}$ (5.14)
@@
$\psi=\ln{\bigg{(}\sqrt{\frac{b_{1}}{\lambda}}\cos^{2}\bigg{(}\pm\sqrt{\frac{\lambda}{2}}(t-\frac{k}{\omega}x)+b_{2}\bigg{)}\bigg{)}}$ (5.15)
@@
Hence,
$\chi(t-\frac{k}{\omega}x)=\exp(\psi)\propto\cos^{2}\bigg{(}\pm\sqrt{\frac{\lambda}{2}}(t-\frac{k}{\omega}x)+b_{2}\bigg{)}\ ,$ (5.16)
with $|\omega|>|k|$, is physically meaningful and is an acceptable solution.
Indeed, we find that it is possible to have exact solutions corresponding to the four-form and circularly polarized, monochromatic GWs with phase velocity exceeding the speed of light. The caveat, however, is that eq.(5.10) being a nonlinear differential equation, deriving exact solutions corresponding to eqs.(5.8) and (5.10) with superposed GW modes with a physically acceptable group velocity is non-trivial.
6 Conclusions
Based on the invariances of the Minkowski tensor and the flat space-time Levi-Civita tensor under proper Lorentz transformations, the present study explores the possibility of extending Einstein’s geometrical formulation of gravitation by including another geometrical field, $\tilde{w}$ a generalisation of the Levi-Civita symbol, in the theory. The geometrodynamical four-form $\tilde{w}$ leads not only to a dynamical torsion, it also generates Chern-Simon (CS) extensions in the 3+1-dimensional space-times. Torsion, among several other interesting implications, is also important in teleparallel gravity theories [4, 22].
Pursuing closely the seminal work of Jackiw and Pi [6] but avoiding the use of an unphysical Lorentz vector, CS coupling between electromagnetic fields and $\tilde{w}$ comes about naturally, bringing about a modification of Einstein-Maxwell equations. Adopting the formalism of [6], a gravitational CS term is also constructed that leads to a modified Cotton tensor. However, when the dynamical equation of the four-form is used in the ensuing Einstein equation, it is shown that it is the standard Cotton tensor that affects the space-time dynamics.